# Supervising 3D Talking Head Avatars with Analysis-by-Audio-Synthesis

### v2

@@ 5.3.1 Quantitative evaluation
Figure 4: Perceptual study of THUNDER. We compare THUNDER-F with methods having both a trainable audio encoder (THUNDER-T w/o m2s) and frozen encoders (FlameFormer-F, THUNDER-F w/o m2s), and GT. The participants prefer THUNDER-F’s lip-sync over that of other models, but does not quite reach the fidelity of GT. Remarkably, the participants also have slight preference for THUNDER-F over GT, which suggests that the application of mesh-to-speech helps THUNDER saturate the quality of GT.
Figure 5: Perceptual study of THUNDER-T. We compare THUNDER-T with methods having both a trainable audio encoders (FlameFormer-T, THUNDER-T w/o m2s) and GT. Models with trainable backbones, achieve good lip-sync than their frozen counterparts, but THUNDER-T is still preferred over the other models.
@@ 5.3.2 Qualitative evaluation
Figure 6: Qualitative comparison. This figure shows the comparison between baselines, our model and GT for selected utterances. The method’s training dataset is listed in parantheses below.
We demonstrate the superiority of our method qualitatively. Fig. 6 shows the comparison of our method with the baselines for selected utterances. Additional qualitative evaluation, can be found in the supplemental video and PDF.
@@ 5.4 Limitations
THUNDER has the following limitations. The absence of teeth and tongue inherently creates ambiguity, making it more difficult to produce high quality audio compared to silent video-to-speech. Further, pseudo-GT is not a perfect reconstruction but has occasional artifacts or inaccuracies. Also, our current mesh-to-speech architecture is fully deterministic and hence is not completely capable of capturing the prediction ambiguities. All of these may result in suboptimal effects of the audio cycle-consistency loss. However, regardless of these shortcomings, our experiments demonstrate the benefits of our method. Employing a large-scale dataset of high quality 3D scans, a face model that models teeth and tongue, and upgrade mesh-to-speech to a stochastic model are likely to boost the benefits of the mesh-to-speech loss. Finally, the inference process of the diffusion model is currently computationally intensive. Future work should address this by employing more efficient solvers such as DPM++ [lu2022dpmpp].
@@ 7.1 Mesh-to-Speech
Per-Dataset Performance Analysis. The performance of both mesh-to-speech and video-to-speech models varies dramatically depending on the dataset. Datasets with limited number of utterances and vocabulary such as RAVDESS or GRID are less challenging, and the models can regress relatively accurately, sometimes to the point of perfect intelligibility. On the more challenging TCD-TIMIT dataset, which contains more complex vocabulary, words are often missed. However, despite the relatively high word error rate (WER) on TCD-TIMIT, the produced sounds, remain very realistic, given the animation. Please see Tab. 3 for per-dataset breakdown and refer to the supplementary video for an audible comparison.
@@
7.2 Talking Heads
@@ 7.2.1 Disentanglement
Ideally, mesh-to-speech should help preserve high-quality lip-sync in the presence of other editing conditions. In other words, the lip-sync should effectively be disentangled from other conditions. To validate this, we evaluate our re-implementations of Media2Face on our test set, passing additional conditioning inputs extracted from emotional images. The conditioning images can be found in Fig. 7. For example results of Media2Face animations, please refer to the supplementary video. The metrics can be found in Tab. 4.
Name LVE $\downarrow$ LIP CCC $\uparrow$ LIP PCC $\uparrow$ DTW $\downarrow$ S-DIV-L $\downarrow$
M2F-F* orig. 0.611 0.522 0.617 0.221 0.0956
M2F-F* w/ m2s orig. 0.551 0.558 0.679 0.212 0.027
M2F-F* angry 1.6 0.307 0.526 0.914 0.151
M2F-F* w/ m2s angry 1.46 0.348 0.59 0.741 0.08
M2F-F* calm 0.91 0.369 0.537 0.297 0.103
M2F-F* w/ m2s calm 0.86 0.459 0.638 0.302 0.0289
M2F-F* disgust 0.928 0.314 0.527 0.316 0.0675
M2F-F* w/ m2s disgust 0.874 0.426 0.627 0.305 0.019
M2F-F* fearful 1.01 0.38 0.618 0.315 0.101
M2F-F* w/ m2s fearful 1.03 0.353 0.627 0.303 0.0295
M2F-F* happy 1.31 0.201 0.394 0.484 0.152
M2F-F* w/ m2s happy 1.11 0.3 0.597 0.269 0.047
M2F-F* sad 0.96 0.308 0.531 0.363 0.108
M2F-F* w/ m2s sad 0.815 0.428 0.609 0.282 0.0371
M2F-F surprised 1.04 0.391 0.576 0.494 0.0759
M2F-F w/ m2s surprised 1.06 0.366 0.602 0.362 0.0278
Table 4: Disentanglement effect. We feed different input conditions to our versions of Media2Face* models and report the lip-sync metrics for both. Orig. denotes an image condition extracted from the video that corresponds to the audio. All other rows are results for denoising with other conditions that come from 7 different emotional images. Media2Face* with THUNDER exhibits superior lip-sync in most cases.
@@ 7.2.2 Extended Ablation and Sensitivity Analysis
Name LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ FDD-U $\downarrow$ FDD-L $\downarrow$
THUNDER w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
@@
THUNDER-F w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
@@
THUNDER-T w/o m2s 0.723 0.428 0.623 0.266 0.02 0.0411 0.0656 0.216 0.0118 0.0811
@@
Table 5: Ablation study and sensitivity analysis. Here we analyze the effect of the mesh-to-speech loss on the talking head avatar training. The top section compares models supervised with mouth-mesh-to-speech (mouth2s), full-mesh-to-speech (face2s) and flame-to-speech (exp2s). While it is already present in the main paper text, we include it here for completeness. The mid and bottom sections analyze the effect of weight $w_{mts}$, with either frozen (-F) or trainable (-T) Wav2Vec2. THUNDER-T models exhibit superior lip-sync performance but it comes at the expense of reduced diversity. Furthermore, for both THUNDER-F and THUNDER-T models, increasing $w_{mts}$, while beneficial for lip animation quality, comes at an increasing expense of of diversity. Please refer to the Sup. Video for qualitative comparison.
Tab. 5 shows the complete ablation of all THUNDER components. We give a comprehensive evaluation of our design decisions. (1) The input space for the mesh-to-speech loss, (2) the weight of the mesh-to-speech loss, and (3) the trainability of the input audio encoder.
On mesh-to-speech input space. While all three modalities are comparable, mouth2s has scores best on the lip-sync metrics, likely thanks to the localized effect on the mouth. Exp2s supervision performs slightly worse on lip-sync but results in higher sample diversity. Finally, applying supervision through the whole face to speech (face2s), while achieving comparable lip-sync scores, scores worse on diversity, specifically S-DIV-U.
On mesh-to-speech strength. Applying mesh-to-speech loss, even weakly, results in improved lip-sync metrics (LVE, CCC, PCC and DTW). The stronger the loss, the more the lip-sync metrics improve. However, increasing weights of the mesh-to-speech loss come at the expense of generation diversity (metrics S-DIV, T-DIV).
On finetuning the audio encoder. Nearly all recent methods finetune the input audio encoder or else they suffer from considerably impaired lip-sync. This results in a certain amount of overfitting to the training subjects’ voices, which manifests as lower diversity scores (S-DIV-U) compared to the models with frozen Wav2Vec (THUNDER-F). In essence, the model’s with trainable audio encoders become less stochastic and more deterministic. Remarkably, the application of the mesh-to-speech loss alleviates the necessity for finetuning Wav2Vec, producing significantly better lip-sync even without finetuning Wav2Vec (THUNDER-T), dramatically reducing the number of training parameters. Applying mesh-to-speech along with finetuning Wav2Vec results in further improvement of lip-sync metrics.

### v3

@@ 5.3.1 Quantitative evaluation
Figure 4: Perceptual study of THUNDER. We compare THUNDER-F (-F for frozen backbone) with methods having both a trainable audio encoder (THUNDER-T w/o m2s) and frozen encoders (FlameFormer-F, THUNDER-F w/o m2s), and GT. The participants prefer THUNDER-F’s lip-sync over that of the other models. Remarkably, the participants also have slight preference for THUNDER-F over GT, which suggests that the application of M2S helps THUNDER saturate the quality of GT.
Perceptual study. We conduct a perceptual study in which we evaluate THUNDER and the effect of mesh-to-speech. Specifically, we compare to the deterministic FlameFormer, THUNDER with and without M2S, and p-GT. The participants are shown two videos side-by-side and asked to rate three aspects of the (lip-sync, dynamism, and realism) animations on a five-point Likert scale (strong/weak preference for A or B or indifferent). We find that participants prefer THUNDER over other methods. Fig. 4 reports the results for THUNDER for lip-sync. For the remaining results of the study and further discussion, see Sec. 12 in Sup. Mat.
@@ 5.3.2 Qualitative evaluation
We refer the reader to the supplemental video for a detail comparisons of the animations, which demonstrates the superiority of our method qualitatively. Fig. 5 shows the comparison of our method with the baselines for selected utterances. Additional qualitative evaluation can be found in Sup. Mat..
@@ 5.4 Limitations
THUNDER has the following limitations. The absence of teeth and tongue inherently creates ambiguity, making it more difficult to produce high quality audio compared to silent video-to-speech. Further, pseudo-GT is not a perfect reconstruction but has occasional artifacts or inaccuracies. Also, our current M2S architecture is fully deterministic and hence is not completely capable of capturing the prediction ambiguities. All of these may result in suboptimal effects of the audio cycle-consistency loss. Regardless of these shortcomings, our experiments demonstrate the benefits of our method. Employing a large-scale dataset of high quality 3D scans, a face model that models teeth and tongue, and upgrade M2S to a stochastic model are likely to boost the benefits of the M2S loss. Finally, the inference process of the diffusion model is currently computationally intensive. Future work should address this by employing more efficient solvers such as DPM++ [54].
Figure 5: Qualitative comparison on THUNDERSET. This figure shows the comparison between baselines, our model and GT for selected utterances. Note that DiffPoseTalk was trained on TFHP. Supplemental PDF and video contain more qualitative comparisons.
@@ 7.1 Mesh-to-Speech
7.1.1 Per-Dataset Performance Analysis.
The performance of both mesh-to-speech and silent-video-to-speech models varies dramatically depending on the dataset. Datasets with limited number of utterances and vocabulary such as RAVDESS or GRID are less challenging, and the models can regress relatively accurately, sometimes to the point of perfect intelligibility. On the more challenging TCD-TIMIT dataset, which contains more complex vocabulary, words are often missed. However, despite the relatively high word error rate (WER) on TCD-TIMIT, the produced sounds remain plausible given the input animation. Please see Tab. 4 for per-dataset breakdown and refer to the supplementary video for an audible comparison.
@@
7.2 Speech-Driven Facial Animation
7.2.1 Experiments on TFHP
Lip-Sync Upper diversity [mm] Lip diversity [mm] Face Dynamic Dev. [mm] Head Pose Chae et. al [8]
Experiment Name Backbone Input LVE [mm] $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW [mm] $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\downarrow$ T-FDD-U $\downarrow$ T-FDD-L $\downarrow$ BA $\times 10^{-3}$ $\uparrow$ DIV $\uparrow$ MTM [ms] $\downarrow$ PLRS $\uparrow$
(1) models trained without head pose THUNDER-F w/o m2s Wav2Vec2 frozen shape, audio 1.21 0.316 0.42 2.37 0.299 0.386 2.7 3.71 0.0739 0.79 n/a n/a 125.141 0.175
THUNDER-F Wav2Vec2 frozen shape, audio 1.05 0.317 0.441 1.93 0.195 0.328 1.5 3.12 0.0627 0.759 n/a n/a 103.185 0.269
DiffPoseTalk HuBERT trainable shape, audio 1.02 0.41 0.526 1.84 0.272 0.383 1.7 2.98 0.0764 0.679 n/a n/a 93.485 0.265
THUNDER-T w/o m2s Wav2Vec2 trainable shape, audio 1.03 0.388 0.525 1.84 0.265 0.364 1.58 2.95 0.0679 0.609 n/a n/a 94.132 0.247
THUNDER-T Wav2Vec2 trainable shape, audio 0.963 0.415 0.55 1.77 0.228 0.341 1.09 2.79 0.0592 0.59 n/a n/a 88.908 0.329
(2) models trained with head pose THUNDER-F w/o m2s Wav2Vec2 frozen shape, audio 1.2 0.285 0.388 2.26 0.258 0.349 2.6 3.58 0.0649 0.76 6.02 3.07 134.403 0.153
THUNDER-F Wav2Vec2 frozen shape, audio 1.07 0.298 0.435 1.95 0.203 0.321 1.5 2.98 0.0707 0.95 6.13 3.78 105.536 0.292
DiffPoseTalk HuBERT trainable shape, audio 1.01 0.414 0.541 1.85 0.247 0.352 1.58 3.02 0.0641 0.618 5.54 2.16 90.005 0.251
THUNDER-T w/o m2s Wav2Vec2 trainable shape, audio 1.12 0.358 0.488 1.86 0.241 0.351 1.87 3.21 0.0597 0.661 6.25 3.49 101.232 0.230
THUNDER-T Wav2Vec2 trainable shape, audio 0.987 0.422 0.55 1.81 0.244 0.341 1.35 2.89 0.0578 0.55 5.77 4.83 89.550 0.320
(3) models trained with head pose and style condition DiffPoseTalk HuBERT trainable audio, shape, style 0.897 0.439 0.555 1.77 0.191 0.319 1.27 2.86 0.0632 0.572 5.57 1.43 88.647 0.256
THUNDER-T w/o m2s Wav2Vec2 trainable audio, shape, style 0.899 0.443 0.545 1.73 0.206 0.31 1.45 2.91 0.0549 0.573 5.72 3.3 94.046 0.261
THUNDER-T Wav2Vec2 trainable audio, shape, style 0.891 0.442 0.567 1.72 0.186 0.31 0.941 2.81 0.0598 0.625 5.98 1.96 89.108 0.327
THUNDER-F w/o m2s Wav2Vec2 frozen audio, shape, style 1.02 0.309 0.386 2.04 0.227 0.323 2.39 3.42 0.0646 0.726 5.85 2.64 134.639 0.124
THUNDER-F Wav2Vec2 frozen audio, shape, style 0.917 0.361 0.465 1.85 0.181 0.312 1.49 2.98 0.0679 0.755 5.92 2.4 103.993 0.280
(4) models trained with perc. loss Chae et al. [8] THUNDER-T w/o m2s, perc. Wav2Vec2 trainable audio, shape 1.117 0.358 0.488 1.864 0.241 0.351 1.87 3.21 0.060 0.661 6.25 3.49 101.232 0.230
THUNDER-T w/o m2s, w/ perc. Wav2Vec2 trainable audio, shape 1.117 0.325 0.485 1.991 0.246 0.344 1.60 3.00 0.064 0.783 5.71 3.99 100.658 0.472
THUNDER-T w/ m2s, w/o perc. Wav2Vec2 trainable audio, shape 0.987 0.422 0.550 1.811 0.244 0.341 1.35 2.89 0.058 0.550 5.77 4.83 89.550 0.320
THUNDER-T w/ m2s, perc. Wav2Vec2 trainable audio, shape 1.079 0.365 0.537 1.869 0.224 0.329 1.11 2.74 0.064 0.777 5.92 2.88 88.862 0.469
THUNDER-F w/o m2s, perc. Wav2Vec2 frozen audio, shape 1.2 0.285 0.388 2.26 0.258 0.349 2.6 3.58 0.0649 0.76 6.02 3.07 134.403 0.153
THUNDER-F w/o m2s, w/ perc. Wav2Vec2 frozen audio, shape 1.2 0.318 0.462 2.27 0.28 0.364 2.07 3.11 0.0652 0.704 5.98 3.66 107.130 0.455
THUNDER-F w/ m2s, w/o perc. Wav2Vec2 frozen audio, shape 1.07 0.298 0.435 1.95 0.203 0.321 1.5 2.98 0.0707 0.95 6.13 3.78 105.536 0.292
THUNDER-F w/ m2s, perc. Wav2Vec2 frozen audio, shape 1.16 0.348 0.498 2.07 0.33 0.375 2.13 3.23 0.0751 0.687 5.9 4.48 99.495 0.479
To verify whether the performance of THUNDER and mesh-to-speech translates to other types of data, we also conduct a comprehensive evaluation on TFHP (the DiffPoseTalk dataset [74]). TFHP was reconstructed from YouTube videos. The dataset contains unscripted audios with unconstrained longer sequences. The reconstructions were obtained using MICA [100] and SPECTRE [29] and look qualitatively very different compared to our EMICA reconstructions of GRID, RAVDESS and TCD-TIMIT datasets. Can THUNDER work also in this setting?
We train both M2S and THUNDER on TFHP and compare it to the official DiffPoseTalk release. To achieve maximum fairness of comparison, we adopt DiffPoseTalk’s number of diffusion steps ($D=500$) and their cosine diffusion schedule in THUNDER. We also use their dataset normalization statistics (i.e. normalize the FLAME coefficients by subtracting mean and dividing by standard deviation). Furthermore, in order to equalize the input conditions, we trained THUNDER with an extra input condition - the FLAME identity shape vector $\boldsymbol{\beta}$, which is also what the original DiffPoseTalk is trained with. We train a number of models to cover all possible scenarios:
(1) Models trained without head pose.
Consistently to the experiments on THUNDERSET, we find that THUNDER models benefit from the M2S loss, which can be observed in improvement on all lip-sync metrics. The improvement is more dramatic in THUNDER-F models, but manifests itself also in the models with trainable audio encoders (THUNDER-F). Similarly to our experiments on THUNDERSET, the application of M2S comes with a small reduction in upper-face diversity. Finally, THUNDER-T outperforms DiffPoseTalk an all lip-sync metrics. The metrics can be seen in Tab. 5, experiment 1.
(2) Models trained with head pose.
Since TFHP comes with head pose annotations that correspond to natural head movement during spontaneous speech, we investigate if THUNDER models can also be trained to produce head pose. When it comes to lip-sync and face diversity, all findings are consistent with that of the experiment without head pose. Additionally, we find that THUNDER is capable of producing head pose with higher beat alignment (BA) and higher diversity (DIV) than DiffPoseTalk. This suggests that the M2S loss does not interfere with head pose generation. See Tab. 5, experiment 2 for more detail. Note that the beat alignment and diversity were computed with the global rotation representation converted into the 6D rotation representation.
(3) Models trained with head pose and speaking style conditioning.
Finally, we train THUNDER models with with the DiffPoseTalk contrastive style features as conditions. The results can be found in See Tab. 5, experiment 3. We find that conditioning with the style vector results in better lip-sync performance, lower face diversity, higher beat alignment and lower pose diversity compared to measurements in See Tab. 5, experiment 2. All of these are consistent with expectations, since passing in a style specification reduces the distribution of possible outcomes. Regardless of the above, the application of M2S has the expected effect - improved lip-sync metrics and lower diversity.
(4) Co-supervising with a perceptual loss.
A recent paper by Chae et al. [8] proposes a new self-supervised speech-mesh representation, in form of a bi-modal masked autoencoder. The architecture consists of two encoders and two decoders (one for the the mesh modality and one for the audio modality). The model is trained with two loss terms - the autoencoder reconstruction losses for both modalities and the InfoNCE loss which brings the two modalities into one coherent features space. This is similar to what CLIP [66] does with images and text. The authors show that their representation can be used in training other speech-driven animation methods. They use their frozen autoencoder in training, passing the input audio and output geometry in, extracting the features for both. Then, they apply the same InfoNCE loss between the two modalities and use it to co-supervise the talking head avatar system. In this experiment, we incorporate this perceptual loss into the THUNDER training. The training loss is the given as:
$\mathcal{L}_{total}=w_{m2s}\mathcal{L}_{m2s}+w_{perc}\mathcal{L}_{perc}+\mathcal{L}_{rec}+\mathcal{L}_{vel},$ (9)
where $w_{perc}\mathcal{L}_{perc}$ is the perceptual term. We follow the authors and the weight accordingly to . For completeness, we also report the new metrics proposed by Chae et al. [8] - mean temporal misalignment (MTM) and Perceptual Lip Readability Score (PLRS) along with the other metrics used in this paper. MTM measures the temporal misalignment using a Derivative Dynamic Time Warping (DDTW). SLRS is the same InfoNCE-based perceptual loss used for training, but computed by a different instance of the same architecture. The authors released two instances - one to be used as the perceptual loss and one to be used for evaluation with SLRS. We utilize both of the speech-mesh autoencoder models released by the authors accordingly. Tab. 5, experiment 4 reports the results of this experiment. In this experiment we investigate the following:
(a) Does training THUNDER with mesh-to-speech result in improvement on metrics proposed by Chae et al. ? Yes, all THUNDER in all settings (across experiments 1-4) improve THUNDER without mesh-to-speech and DiffPoseTalk on both MTM and PLRS. The improvement on MTM is expected, since MTM leverages DDTW, which is similar to the DTW metric [84] we use. Furthermore, improvement on PLRS indicates that the mesh-to-speech loss results in a perceptual improvement in quality.
(b) Does training our diffusion-based speech-driven animation method be trained with this perceptual loss? Yes, THUNDER models trained without mesh-to-speech but with the perceptual loss improve considerably on PLRS. This is expected since PLRS is computed by a model with the same architecture as the network which provides the perceptual loss function. We also observe improvement on MTM. Furthermore, we observe improvement on our lip-sync related metrics in models with the frozen backbone (-F). Models with trainable backbones (-T) do not seem to improve on LVE, CCC, PCC or DTW when the perceptual loss is applied.
(c) Can THUNDER be trained with both mesh-to-speech and the perceptual loss. Yes. Both losses can be applied together. The mesh-to-speech loss has a strong impact on LVE, CCC, PCC, DTW and MTM, while the perceptual loss strongly affects PLRS, MTM and with -F models also improves CCC and PCC. THUNDER-F with both mesh-to-speech and the perceptual loss appears to get the best of both worlds, while THUNDER-T with mesh-to-speech only is competetive with THUNDER-T with both losses. In any case, applying the mesh-to-speech is beneficial.
@@ 7.2.2 Disentanglement
Ideally, M2S should help preserve high-quality lip-sync in the presence of other editing conditions. In other words, the lip-sync should effectively be disentangled from other conditions. To validate this, we evaluate our re-implementation of Media2Face (THUNDERSET-trained) on the THUNDERSET test set, passing additional conditioning inputs extracted from emotional images from unrelated sequences. The conditioning images can be found in Fig. 6. For example results of Media2Face animations, please refer to the supplementary video. We find that training Media2Face with mesh-to-speech improves lip-sync. The metrics can be found in Tab. 6.
Name LVE [cm] $\downarrow$ LIP CCC $\uparrow$ LIP PCC $\uparrow$ DTW [cm] $\downarrow$ S-DIV-L [cm] $\downarrow$ MTM [ms] $\downarrow$ PLRS $\uparrow$
M2F-F* orig. 0.611 0.522 0.617 0.221 0.0956 41.0 0.27
M2F-F* w/ m2s orig. 0.551 0.558 0.679 0.212 0.027 57.5 0.283
M2F-F* angry 1.6 0.307 0.526 0.914 0.151 88.8 0.173
M2F-F* w/ m2s angry 1.46 0.348 0.59 0.741 0.08 61.7 0.176
M2F-F* calm 0.91 0.369 0.537 0.297 0.103 75.5 0.257
M2F-F* w/ m2s calm 0.86 0.459 0.638 0.302 0.0289 62.5 0.27
M2F-F* disgust 0.928 0.314 0.527 0.316 0.0675 77.9 0.249
M2F-F* w/ m2s disgust 0.874 0.426 0.627 0.305 0.019 66.6 0.296
M2F-F* fearful 1.01 0.38 0.618 0.315 0.101 60.8 0.263
M2F-F* w/ m2s fearful 1.03 0.353 0.627 0.303 0.0295 56.4 0.24
M2F-F* happy 1.31 0.201 0.394 0.484 0.152 82.0 0.209
M2F-F* w/ m2s happy 1.11 0.3 0.597 0.269 0.047 66.6 0.265
M2F-F* sad 0.96 0.308 0.531 0.363 0.108 77.3 0.257
M2F-F* w/ m2s sad 0.815 0.428 0.609 0.282 0.0371 56.9 0.287
M2F-F* surprised 1.04 0.391 0.576 0.494 0.0759 63.5 0.26
M2F-F* w/ m2s surprised 1.06 0.366 0.602 0.362 0.0278 53.7 0.254
Table 6: Disentanglement effect. We feed different input conditions to our versions of Media2Face* models and report the lip-sync metrics for both. Orig. denotes an image condition extracted from the video that corresponds to the audio. All other rows are results for denoising with other conditions that come from 7 different emotional images. Media2Face* with THUNDER exhibits superior lip-sync in most cases. Please note that since there is not Ground Truth available for any other setting except for passing the original image (ie. M2F orig.), the metrics are still computed against the original GT. That said, since PLRS is a perceptual score, it does not require GT. However, the consistent improvement in the lip metrics across all condition is still indicative of superior lip-sync.
@@ 7.2.3 Extended Ablation and Sensitivity Analysis
Lip-Sync Upper-face diversity Lip diversity Face Dynamic Dev.
Experiment Name LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\downarrow$ FDD-U $\downarrow$ FDD-L $\downarrow$
(1) THUNDER and mesh-to-speech input space THUNDER w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
@@
(2) THUNDER-F mesh-to-speech sensitivity analysis THUNDER-F w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
@@
(3) THUNDER-T mesh-to-speech sensitivity analysis THUNDER-T w/o m2s 0.723 0.428 0.623 0.266 0.02 0.0411 0.0656 0.216 0.0118 0.0811
@@
Table 7: Ablation study and sensitivity analysis. Here we analyze the effect of the mesh-to-speech loss on the talking head avatar training. This experiment is conducted on THUNDERSET. The top section compares models supervised with mouth-mesh-to-speech (mouth2s), full-mesh-to-speech (face2s) and flame-to-speech (exp2s). While it is already present in the main paper text, we include it here for completeness. The mid and bottom sections analyze the effect of weight $w_{m2s}$, with either frozen (-F) or trainable (-T) Wav2Vec2. THUNDER-T models exhibit superior lip-sync performance but it comes at the expense of reduced diversity. Furthermore, for both THUNDER-F and THUNDER-T models, increasing $w_{m2s}$, while beneficial for lip animation quality, comes at an increasing expense of of diversity. Please refer to the Sup. Video for qualitative comparison.
Tab. 7 shows the complete ablation and sensitivity analysis of all THUNDER components on THUNDERSET. Here we give a comprehensive evaluation of our design decisions. (1) The input space for the mesh-to-speech loss, (2) the weight of the mesh-to-speech loss, and (3) the effect if finetuning of the input audio encoder.
(1) On mesh-to-speech input space. While all three modalities are comparable, mouth2s has scores best on the lip-sync metrics, likely thanks to the localized effect on the mouth. Exp2s supervision performs slightly worse on lip-sync but results in higher sample diversity. Finally, applying supervision through the whole face to speech (face2s), while achieving comparable lip-sync scores, scores worse on diversity, specifically S-DIV-U. The metric can be found in Tab. 7, experiment 1.
(2) On mesh-to-speech strength. Applying mesh-to-speech loss, even weakly, results in improved lip-sync metrics (LVE, CCC, PCC and DTW). The stronger the loss, the more the lip-sync metrics improve. However, increasing weights of the mesh-to-speech loss come at the expense of generation diversity (metrics S-DIV, T-DIV). See Tab. 7, experiment 2.
(3) On finetuning the audio encoder. Nearly all recent methods finetune the input audio encoder or else they suffer from considerably impaired lip-sync. This results in a certain amount of overfitting to the training subjects’ voices, which manifests as lower diversity scores (S-DIV-U) compared to the models with frozen Wav2Vec (THUNDER-F). In essence, the model’s with trainable audio encoders become less stochastic and more deterministic. Remarkably, the application of the mesh-to-speech loss alleviates the necessity for finetuning Wav2Vec, producing significantly better lip-sync even without finetuning Wav2Vec (THUNDER-T), dramatically reducing the number of training parameters. Applying mesh-to-speech along with finetuning Wav2Vec results in further improvement of lip-sync metrics. Refer to Tab. 7, experiment 3.